A wall is hit elastically and normally by n balls per second, all the balls have the same mass m and are moving with the same velocity u. the force exerted by the balls on the wall is_______

Grade:12

2 Answers

Saurabh Koranglekar
3 years ago

Since the ball rebounds with same speed but in opposite direction.
So, change in linear momentum of one ballΔP=Pf​−Pi​=mu−(−mu)=2mu
Number of ball hitting the surface per second isn
So, total change in momentum in one secondΔPT​=n×2mu
So, force acting on the surfaceF=2mnu
Vikas TU
3 years ago
When a ball hits normally on surface elastically then it rebounds with same velocity.
F = m * dv/dt
F = nM * (u - (-u)) / t
F = 2nM / t
Given that n/t = N
∴ F = 2NMu
Force on the surface is 2NMu